OpenStudy (anonymous):

tan a = 24/7 and a is in QIII. What is cos(a – pi)? 7/25 -7/25 -24/7 24/7

OpenStudy (anonymous):

Is this correct? invtan(24/7)= 1.2870022175865687736056184574346. This is in quarter I so I add pie to it, but then subtract it in cos(a-pi). I then cosine it to get .28, or 7/25. Did i do it correctly?

OpenStudy (anonymous):

Lets start with process of elimination first then start solving should we

OpenStudy (anonymous):

you in q3 tangent is positive right?

OpenStudy (campbell_st):

draw a triangle  find x using pythagoras

OpenStudy (campbell_st):

next find sin(a) and cos(a) after finding x then use cos(A - B) = cos(A)cos(B) + sin(A)sin(B)

OpenStudy (anonymous):

Magnitude is 25

OpenStudy (campbell_st):

that will give you the value of cos(a - pi)

OpenStudy (anonymous):

You gave me a sum and difference identity, i do not see an A or B

OpenStudy (anonymous):

Ahh

OpenStudy (campbell_st):

you use the identity... let A = a and B = pi does that make sense

OpenStudy (anonymous):

-cos(a)

OpenStudy (campbell_st):

and you should know what \[\cos(\pi) = ?.... and..... \sin(\pi) = \]

OpenStudy (campbell_st):

oops the substitution should be \[\cos(a - \pi) = \cos(a)\cos(\pi) + \sin(a)\sin(\pi)\]

OpenStudy (anonymous):

cos(pi)=-1 sin(pi)=0 so it should reduce to what i just said of -cos(a) right?

OpenStudy (campbell_st):

ok... from the triangle in the 3rd quadrant, what is the value of cos(a)..?

OpenStudy (campbell_st):

that's correct... but you need to find cos(a) in this quadrant...

OpenStudy (anonymous):

-0.28=cos(a)

OpenStudy (campbell_st):

if you look at the triangle are you happy that cos(a) = -7/25 sin(a) = -24/25

OpenStudy (anonymous):

im happy? ok lol yep

OpenStudy (campbell_st):

\[\cos(a - \pi) = \cos(a)\cos(\pi) + \sin(a)\sin(\pi)\] which is \[\cos(a - \pi) = \frac{-7}{25} \times - 1 + \frac{-24}{25} \times 0\] which will give you an answer... from your list of choices

OpenStudy (anonymous):

which gives me 7/24 im sure theres a reason but why is doing it like that better than cos((invtan(24/7)+pi)= 7/24

OpenStudy (campbell_st):

because, looking at how you managed you're solution, it wasn't an exact value. The question gave answer choices that are exact values.... for me the question is asking you to use your trigonometry skills. Show what you know. my method shows a lot of understanding...

OpenStudy (anonymous):

ok